# The Fate of Entanglement

### v2

@@ I.1 Logarithmic negativity
$\mathcal{E}=\log\big{\lVert}\rho^{T_{1}}\!\big{\rVert}$ (S1)
@@ I.2 Geometric entanglement
$\mathcal{D}=\min_{\rho_{\rm sep}}\sqrt{\operatorname{Tr}(\rho-\rho_{\rm sep})^{2}}.$ (S2)
Numerically, we choose a number of product states (up to 7) in the convex combination, and optimize the distance function over these separable states.
I.3 Genuine 3-party entanglement criterion
@@
$W=|\rho_{23}|+|\rho_{25}|+|\rho_{35}|-\sqrt{\rho_{11}\rho_{44}}-\sqrt{\rho_{11}\rho_{66}}-\sqrt{\rho_{11}\rho_{77}}-\frac{1}{2}(\rho_{22}+\rho_{33}+\rho_{55}).$ (S3)
@@
II B. Test of the full-rank hypothesis
@@
We test the full-rank hypothesis (FRH) for the icosahedral molecule at $h=3$ for three adjacent sites. In Fig. S1 we plot the minimal eigenvalues $\lambda_{\rm min}$ of the 3-spin reduced density matrices for each eigenstate of the Hamiltonian, labelled by the corresponding energy $E$. All the minimal eigenvalues are strictly positive, which indicates that the FRH is satisfied. The smallest $\lambda_{\rm min}$ occurs in the groundstate.
@@
Figure S1: Test of the full-rank hypothesis. Minimal eigenvalue $\lambda_{\rm min}$ of the reduced density matrices for three adjacent spins in the icosahedral molecule at $h=3$ for the whole spectrum, labelled by the corresponding energy $E$.
III C. Phase diagram of the molecular quantum magnet
@@
We report the zero-temperature connected correlation functions of adjacent spins in the anti-ferromagnetic Ising model on the icosahedral molecule as a function of the transverse field $h$, see Fig. S2. By symmetry we have $\langle\sigma^{x,y}\rangle=0$. The point $h=3$ corresponds to a generic point of the phase diagram.
@@
Figure S2: Spin correlations in the icosahedral molecule. Various spin correlation functions versus the transverse field in the icosahedral molecule at zero temperature.
IV D. Details on the quench protocol
@@
In the quench protocol, we initialise the system in a pure product state of up and down spins, and let it evolve under the Ising Hamiltonian discussed above with $h=3$. We illustrate the initial state in Fig. S3. This is a planar representation of the icosahedral molecule, where blue sites are initialised in a spin-up state, whereas orange one are initialised in a spin-down state. The coloured face represents the 3-spin subsystem for which we compute $\mathcal{D}$ and $W$ during the time evolution.
@@
Figure S3: Initial state in the quench protocol. Blue sites are initialised in a spin-up state, whereas orange one are initialised in a spin-down state. We focus on the coloured face during the time evolution.

### v4

@@ A.1 Logarithmic negativity
$\mathcal{E}=\log\big{\lVert}\rho^{T_{1}}\!\big{\rVert}$ (8)
@@ A.2 Geometric entanglement
$\mathcal{D}=\min_{\rho_{\rm sep}}\sqrt{\operatorname{Tr}(\rho-\rho_{\rm sep})^{2}}.$ (9)
$\mathcal{D}_{b}$ is analogously defined as the distance to the biseparable set. For a small number of spins, we can perform the optimization to get good upper bounds. We have used $\mathcal{D}_{(b)}$ for 3 spins, while the Gilbert algorithm was used to obtain $\mathcal{D}$ for 4 spins owing to the larger parameter space. In performing the optimization for $\mathcal{D}$ in the case of 3 spins, we search over the set $\sum_{i=1}^{u}K_{i}$ with a fixed number of components $u$, where $K_{i}=\rho_{1}^{i}\otimes\rho_{2}^{i}\otimes\rho_{3}^{i}$ is a general product state (not necessarily pure). We begin with a low $u$, such as $u_{1}=5$, find the optimum by a brute force search (we employed the global routine ‘fmincon’ in Matlab) using many initial points. Once the closest separable state is found, we used as the starting guess for the search with a larger number of components, $u_{2}=u_{1}+1$. Once the variation of $\mathcal{D}$ between successive $u$-values becomes less than a given threshold, we stop the process. We have found that we can get excellent upper bounds by going up to $u=12$. The values obtained were cross-checked with the Gilbert approach. An analogous approach yields $\mathcal{D}_{b}$, but now the product states are replaced by $\rho_{1}\otimes\rho_{23}+\sigma_{2}\otimes\sigma_{13}+\varrho_{3}\otimes\varrho_{12}$, where the subscript denotes the spins of the corresponding physical density matrix.
The separability/biseparability certification through the GTC is obtained via an optimization over the same $u$-component set, as described in [44].
A.3 Genuine 3-party entanglement criterion
@@
$W=|\rho_{23}|+|\rho_{25}|+|\rho_{35}|-\sqrt{\rho_{11}\rho_{44}}-\sqrt{\rho_{11}\rho_{66}}-\sqrt{\rho_{11}\rho_{77}}-\frac{1}{2}(\rho_{22}+\rho_{33}+\rho_{55}).$ (10)
@@
Appendix B Test of the full-rank hypothesis
@@
We test the full-rank hypothesis (FRH) for the icosahedral molecule at a representative value of the transverse field, $h=3$, for three adjacent sites, as well as for four adjacent sites in the spin chain. In Fig. 5 we plot the minimal eigenvalues $\lambda_{\rm min}$ of the 3- and 4-spin reduced density matrices for each eigenstate of the Hamiltonian, labeled by the corresponding energy $E$. All the minimal eigenvalues are strictly positive, which indicates that the FRH is satisfied. For the icosahedral molecule, the groundstate has the smallest $\lambda_{\rm min}$, whereas it is the second smallest in the spin chain.
@@
Figure 5: Test of the full-rank hypothesis. Minimal eigenvalue $\lambda_{\rm min}$ of the reduced density matrices for three adjacent spins in the icosahedral molecule at $h=3$ (top), and for four adjacent spins in the spin chain (bottom) for the whole spectrum, labeled by the corresponding energy $E$. We note that the groundstate has the smallest $\lambda_{\rm min}$ in the top panel, while it is the second smallest in the bottom one.
Appendix C Phase diagram of the molecular quantum magnet
@@
We report the zero-temperature connected correlation functions of adjacent spins in the anti-ferromagnetic Ising model on the icosahedral molecule as a function of the transverse field $h$, see Fig. 6. By symmetry we have $\langle\sigma^{x,y}\rangle=0$. The point $h=3$ corresponds to a generic point of the phase diagram.
@@
Figure 6: Spin correlations in the icosahedral molecule. Various spin correlation functions versus the transverse field in the icosahedral molecule at zero temperature.
Appendix D Quench protocols
@@
In the quench protocols Ico-1 and Ico-2, we initialize the system in a pure product state of up and down spins in the $\sigma^{z}$ basis, and let it evolve under the icosahedron Ising Hamiltonian discussed above with $h=3$. We illustrate the initial states in Fig. 7. This is a planar representation of the icosahedral molecule, where blue sites are initialized in a spin-up state, whereas orange one are initialized in a spin-down state. The colored face represents the 3-spin subsystem for which we compute various metrics during the time evolution. For the Ico-3 protocol, the initial state is the groundstate of the Ising Hamiltonian with $h=0.04$ and it evolves under the Hamiltonian with $h=3$. The time-evolution of the geometric entanglement $\mathcal{D}$ for both Ico-2 and Ico-3 quenches is represented in Fig. 8. We have further studied the 4-spin subregion made of two adjacent triangular faces, and were able to certify full separability for $t\geqslant 4$ (Ico-2) and $t\geqslant 1.5$ (Ico-3).
@@
Figure 7: Initial states in the quench protocols. Blue sites are initialized in a spin-up state, whereas orange ones are initialized in a spin-down state. We study the colored face during the time evolution. The left panel represents the initial stat for the Ico-1 protocol, and the left panel pertains to the Ico-2 protocol.
Figure 8: Icosahedron quench protocols Ico-2 and Ico-3. The geometric entanglement $\mathcal{D}$ and $\mathcal{D}_{b}$ for the quench protocols Ico-2 and Ico-3. The open shapes indicate that the 3-spin state is rigorously certified to be fully (bi)separable.
Appendix E Genuine multipartite entanglement for fermions
In this appendix, we show that fermionic separable states are surrounded in their close vicinity by regions of fermionic biseparability, implying the sudden death of genuine multipartite entanglement (GME) in fermionic systems.
E.1 Three parties
Let us first consider the case of $m=3$ fermionic modes. A fermionic separable state has the form
$\rho_{\rm sep}^{F}=\sum_{k}p_{k}\ \rho_{1}^{(k)}\otimes\rho_{2}^{(k)}\otimes\rho_{3}^{(k)}$ (11)
with $p_{k}\geqslant 0$ and $\sum_{k}p_{k}=1$. The 1-mode states $\rho_{j}^{(k)}$ are normalized, Hermitian and positive semi-definite operators with even local fermion parity: $(-1)^{n_{j}}\rho_{j}^{(k)}(-1)^{n_{j}}=\rho_{j}^{(k)}$, where the number operator $n_{j}=c_{j}^{\dagger}c_{j}$ is defined in terms of the mode annihilation operator $c_{j}$. The generic form for such states is
$\rho_{j}^{(k)}=\frac{1}{2}\mathbb{I}+a_{j}^{(k)}\sigma^{z},\qquad-\frac{1}{2}\leqslant a_{j}^{(k)}\leqslant\frac{1}{2}.$ (12)
We take $\rho_{\rm sep}^{F}$ to lie in the interior of the fermion separable set, for otherwise entangled states would lie in the immediate vicinity of the state. By inclusion, $\rho_{\rm sep}^{F}$ also lies in the interior of the bosonic separable continent. Using convexity of separable states, we are thus guaranteed that there exists a region surrounding $\rho_{\rm sep}^{F}$ where states have a (bosonic) separable form. We parametrize them by
$\rho(\epsilon)=\sum_{k}p_{k}\ (\rho_{1}^{(k)}+\epsilon\ \omega_{1}^{(k)})\otimes(\rho_{2}^{(k)}+\epsilon\ \omega_{2}^{(k)})\otimes(\rho_{3}^{(k)}+\epsilon\ \omega_{3}^{(k)})$ (13a)
with
$\omega_{j}^{(k)}=b_{j}^{(k)}\sigma^{x}+d_{j}^{(k)}\sigma^{y},$ (13b)
where $\epsilon\geqslant 0$ is a positive but small parameter. These states satisfy $\rho(0)=\rho_{\rm sep}^{F}$. In principle, one could consider a more generic perturbation where $\omega_{j}^{(k)}$ also contains a $\sigma^{z}$ component. However, this would be equivalent to considering a state $\tilde{\rho}(\epsilon)$ as in (13a) but centered around another fermionic separable state $\tilde{\rho}_{\rm sep}^{F}$. We thus consider the case where $\omega_{j}^{(k)}$ only contains terms breaking the local fermion parity. We stress that the 1-mode states in (13a) break local fermion parity, and therefore $\rho(\epsilon)$ corresponds to an entangled fermionic state.
The constants $b_{j}^{(k)}$ and $d_{j}^{(k)}$ in (13b) are not entirely free however. They must be chosen such that (i) the individual states in the convex combination (13a) are positive semi-definite (they are normalized and Hermitian by definition), and (ii) the total state commutes with the total fermion parity, $(-1)^{n}$, where $n=\sum_{j}n_{j}$. These conditions translate to, respectively,
$(a_{j}^{(k)})^{2}+\epsilon^{2}(b_{j}^{(k)})^{2}+\epsilon^{2}(d_{j}^{(k)})^{2}\leqslant\frac{1}{4}$ (14a)
and
$\sum_{k}p_{k}\epsilon\!\left(\omega_{1}^{(k)}\otimes\rho_{2}^{(k)}\otimes\rho_{3}^{(k)}+\rho_{1}^{(k)}\otimes\omega_{2}^{(k)}\otimes\rho_{3}^{(k)}+\rho_{1}^{(k)}\otimes\rho_{2}^{(k)}\otimes\omega_{3}^{(k)}\right)\\ +\sum_{k}p_{k}\epsilon^{3}\omega_{1}^{(k)}\otimes\omega_{2}^{(k)}\otimes\omega_{3}^{(k)}=0\,.$ (14b)
The first condition (14a) can always be satisfied for small enough $\epsilon$, unless the local fermionic state in $\rho_{\rm sep}^{F}$ is $\rho_{j}^{(k)}=\tfrac{1}{2}(\mathbb{I}\pm\sigma^{z})$. In those situations, no parity-breaking perturbation on $\rho_{j}^{(k)}$ is allowed and $\omega_{j}^{(k)}=0$.
Because all the odd-parity terms have to cancel, see (14b), we recast $\rho(\epsilon)$ as
$\rho(\epsilon)=\rho_{\rm sep}^{F}+\epsilon^{2}\sum_{k}p_{k}\left(\rho_{1}^{(k)}\otimes\omega_{2}^{(k)}\otimes\omega_{3}^{(k)}+\omega_{1}^{(k)}\otimes\rho_{2}^{(k)}\otimes\omega_{3}^{(k)}+\omega_{1}^{(k)}\otimes\omega_{2}^{(k)}\otimes\rho_{3}^{(k)}\right).$ (15)
For each $k$ in the sum for $\rho(\epsilon)$, we recast the corresponding term (dropping the indices and tensor product symbols for readability) as
$\rho\rho\rho+\epsilon^{2}(\rho\omega\omega+\omega\rho\omega+\omega\omega\rho)=\frac{1}{3}\rho(\rho\rho+3\epsilon^{2}\omega\omega)+\frac{1}{3}(\rho\rho\rho+3\epsilon^{2}\omega\rho\omega)+\frac{1}{3}(\rho\rho+3\epsilon^{2}\omega\omega)\rho\,.$ (16)
We thus have
$\rho(\epsilon)=\sum_{k}\frac{p_{k}}{3}\left(\rho_{1}^{(k)}\otimes\rho_{23}^{(k)}(\epsilon)+\rho_{13}^{(k)}(\epsilon)\otimes\rho_{2}^{(k)}+\rho_{12}^{(k)}(\epsilon)\otimes\rho_{3}^{(k)}\right)$ (17a)
with
$\rho_{ij}^{(k)}(\epsilon)=\rho_{i}^{(k)}\otimes\rho_{j}^{(k)}+3\epsilon^{2}\omega_{i}^{(k)}\otimes\omega_{j}^{(k)}.$ (17b)
For $0\leqslant\epsilon\leqslant\epsilon^{*}$ and $\epsilon^{*}>0$ small enough, this state is guaranteed to be positive semi-definite for all $i,j,k$. The only potential obstacle would be if one $\rho_{j}^{(k)}$ were of the form $\tfrac{1}{2}(\mathbb{I}\pm\sigma^{z})$ (with a vanishing eigenvalue), but in that case the corresponding perturbation would have to vanish, $\omega_{j}^{(k)}=0$, as explained above. Moreover, it is direct to show that condition (14a) is fulfilled. We thus conclude that for $0\leqslant\epsilon\leqslant\epsilon^{*}$, the state $\rho(\epsilon)$ has a biseparable form (2), where each state involved in the convex combination is normalized, Hermitian, positive semi-definite and commutes with its local fermion parity. It it thus fermionic biseparable.
E.2 Four parties
For $m=4$, the argument is similar. The even terms in the sum for $\rho(\epsilon)$, following the notation of (16), have the form
$\rho\rho\rho\rho+\epsilon^{4}\omega\omega\omega\omega+\epsilon^{2}(\rho\rho\omega\omega+\textrm{perm.})=\\ \frac{1}{7}(\rho\rho\rho\rho+7\epsilon^{4}\omega\omega\omega\omega)+\left\{\frac{1}{7}(\rho\rho\rho\rho+7\epsilon^{2}\rho\rho\omega\omega)+\textrm{perm.}\right\}.$ (18)
We recast the terms as
$\begin{split}\rho\rho\rho\rho+7\epsilon^{4}\omega\omega\omega\omega&=\frac{1}{2}(\rho\rho+\sqrt{7}\epsilon^{2}\omega\omega)(\rho\rho+\sqrt{7}\epsilon^{2}\omega\omega)+\frac{1}{2}(\rho\rho-\sqrt{7}\epsilon^{2}\omega\omega)(\rho\rho-\sqrt{7}\epsilon^{2}\omega\omega),\\[8.5359pt] \rho\rho\rho\rho+7\epsilon^{2}\rho\rho\omega\omega&=\rho\rho(\rho\rho+7\epsilon^{2}\omega\omega),\end{split}$ (19)
and similarly for the other five permutations involving two $\rho$ and two $\omega$. As for the case $m=3$, there are small but non-vanishing values of $\epsilon$ for which all states above are positive semi-definite. Since they commute with their local parity, it means that the total state $\rho(\epsilon)$ is fermionic biseparable. The argument readily generalises to arbitrary values of $m\geqslant 3$.